OpenStudy (anonymous):

could you help me with Solve the system by substitution. y = 2x 3x + 2y = 21 A.x = 3, y = 6 B.x = 6, y = 3 C.x = 3, y = –6 D.x = –6, y = 3

OpenStudy (anonymous):

sure thing. Its fairly straight forward, as one of the equations gives you that y is 2x now substitute that into the second equation to get 3x + 2(2x) = 21 this simplifies to 3x+4x=21 7x=21 7x/7=21/7 x=3 now substitute this back into your first equation y=2x=2(3)=6 x=3 y=6

OpenStudy (anonymous):

ooooooo....thanks!

OpenStudy (anonymous):

the problem goes on to say... Why is (6, –2) not a solution to the system of equations given? 2x-3y= 18 5x=3y=3

OpenStudy (anonymous):

is that second equation supposed to have 2 equals signs?